OpenStudy (anonymous):

Will fan and medal. Stuck on this question heh 6 1/3 + k + 3 5/6 + 5 1/2 = 26 1/6

OpenStudy (tkhunny):

Have you considered:...? 1) Changing mixed numbers to improper fractions 2) Finding a common denominator

OpenStudy (anonymous):

First, get everything to a common denominator: 6 1/3 becomes 6 2/6 5 1/2 becomes 5 3/6 Now we have: 6 2/6 +k + 3 5/6 + 5 3/6 = 26 1/6 Next, Isolate k k = 26 1/6 - (6 2/6 + 3 5/6 + 5 3/6) Simplify k = 26 1/6 - (15 4/6) Simplify k = 10 3/6 Simplify k = 10 1/2 or 10.5

OpenStudy (tkhunny):

Why didn't 6 become 36/6 or k become k/6? "Get everything" doesn't seem to be where you went.

OpenStudy (anonymous):

I changed the mixed numbers into improper fractions, aaaaand I got it. Thank you c: